# NaRCan: Natural Refined Canonical Image with Integration of Diffusion Prior for Video Editing

### v1

@@ Abstract
We propose a video editing framework, NaRCan, which integrates a hybrid deformation field and diffusion prior to generate high-quality natural canonical images to represent the input video. Our approach utilizes homography to model global motion and employs multi-layer perceptrons (MLPs) to capture local residual deformations, enhancing the model’s ability to handle complex video dynamics. By introducing a diffusion prior from the early stages of training, our model ensures that the generated images retain a high-quality natural appearance, making the produced canonical images suitable for various downstream tasks in video editing, a capability not achieved by current canonical-based methods. Furthermore, we incorporate low-rank adaptation (LoRA) fine-tuning and introduce a noise and diffusion prior update scheduling technique that accelerates the training process by 14 times. Extensive experimental results show that our method outperforms existing approaches in various video editing tasks and produces coherent and high-quality edited video sequences. See our project page for video results: koi953215.github.io/NaRCan_page.
@@ 1 Introduction
Video editing has always been a fascinating research area. For example, style transfer transforms the original video into a completely new style, enriching the viewing experience. Other tasks include dynamic segmentation and handwriting, which all demonstrate the broad application value of video editing across various fields. Currently, diffusion model technology is becoming increasingly mature and is known for its powerful generative capabilities and frequent use in video editing. However, in video-to-video tasks, maintaining temporal consistency is crucial. Generally, diffusion models that are not specially processed cannot generate video sequences with sufficient temporal consistency. This has led to numerous papers addressing this issue, attempting to make diffusion models produce high-quality video sequences while ensuring sequence consistency.
Yet, even if the problem of temporal consistency is solved, diffusion-based methods [16, 4, 19, 8] still face challenges in adapting to certain downstream tasks, such as handwriting. This is where canonical-based methods come into play. Canonical-based methods [3, 1, 5, 44] generate a canonical image representing all the video information, meaning that editing this single image equates to editing the entire video. Thus, these methods can easily adapt to various video editing tasks.
It is important to note that this is only true if the canonical image is a high-quality, natural image. If the canonical image is not natural, it loses any editing value. We observed that existing canonical-based methods do not incorporate any constraints to ensure the canonical image is of high quality and natural. To address this issue, we propose NaRCan, a novel hybrid deformation field network architecture that integrates diffusion priors (Figure 1) into our training pipeline (Figure 2) to ensure the generation of high-quality, natural canonical images across various scenarios. Through a series of experiments, we demonstrate that NaRCan outperforms existing video editing methods.
@@ 2 Related Work
Some works utilize GAN inversion [64, 76, 59, 72, 47, 26] to edit images or videos. Today, numerous generative models exist for editing images. Some methods, such as GLIDE [43], DALL-E [54, 53], stable diffusion [55], and Imagen [57], are trained on millions of images, resulting in incredible generative abilities. Other methods like SDEdit [38], ControlNet [73], and LDM [2] use conditions to achieve better editing results. Instruction-based video editing methods like InstructPix2Pix [3] and another work [10] often yield sub-optimal results for different editing operations. Techniques like LoRA [18] can assist in fine-tuning to find better weights for editing. Additionally, many zero-shot diffusion-based methods [70, 16, 74, 4, 8, 19, 27] do not require model training but still need constraints to maintain temporal consistency. To address temporal consistency concerns, most methods like Tune-A-Video [68], Text2Video-Zero [27], FateZero [51], and Vid2Vid-Zero [65] incorporate cross-attention mechanisms. Some works propose training for video editing, such as Imagen Video [17] and Make-A-Video [58], but these require large datasets and significant computational resources. Unlike these methods, MeDM [13] uses a flow-coding algorithm to solve this problem. Canonical-based design does not require another mechanism to maintain temporal consistency, however. Once the canonical image is edited, the changes can be propagated to every frame using a deformation field.
@@ 3 Method
In this section, we first introduce our hybrid deformation modeling by combining homography and deformation MLP in Section 3.1. Next, we elaborate on how we integrate the diffusion prior from a LoRA fine-tuned latent diffusion model to ensure the naturalness of our canonical image representation in Section 3.2. Finally, we provide an additional way to improve the quality of our video representation by separating multiple canonical images and describing the necessary changes for downstream tasks in Section 3.3. Figure 2 shows our proposed framework.
@@ 3.1 Hybrid Deformation Field for Deformation Modeling
supplemented with a TVFlow regularization term to prevent overfitting by limiting its expressive capacity. However, this regularization term restricts the model’s ability to express itself without providing additional insights. Therefore, we propose a hybrid deformation field architecture composed of a trainable homography matrix $H(u,v,t)$ and residual deformation MLP:
@@
Unlike TVFlow, homography provides global displacement information as heuristic information for the subsequent residual deformation MLP. This enables the residual deformation MLP to learn and express the deformation field more accurately and effectively (Figure 9(a)).
@@ 4.2 Evaluation
Method Venue Warping error $\downarrow$ Interpolation error $\downarrow$
Hashing-nvd [5] ICCV 2023 0.594 9.204
CoDeF [44] CVPR 2024 0.486 8.721
MeDM [13] AAAI 2024 0.367 9.941
Ours - 0.364 8.365
@@
We run our method CoDeF [44], Hashing-nvd [5], CCEdit  [15], and MeDM [13] on the DAVIS [48] and BalanceCC Benchmark that CCEdit proposed for evaluating the results of video editing. In Figure 5, we show the comparison of visual results. BalanceCC Benchmark provided a unified text prompt for each scene, and we ensured that all videos from the BalanceCC Benchmark were kept within 100 frames. We utilized ControlNet [73] to edit the video using the unified text prompt provided by BalanceCC Benchmark for a fair comparison. We executed MeDM and CCEdit within their provided environment settings and pre-trained models. Moreover, Hashing-nvd is a video decomposition work that outputs two images representing foreground and background. To maintain the consistent style of these two images for video editing, We utilize the grid trick proposed by RAVE [24] to tackle this problem. Finally, there are also some video-editing results on DAVIS [48], and the corresponding text prompts originated from the BalanceCC Benchmark. In Table 1, We choose warping error [29, 75] and interpolation error [31] as our metrics of temporal consistency.
@@ 5 Conclusion
Our method relies on LoRA [18] fine-tuning to enhance the diffusion model’s ability to represent the current scene. However, LoRA fine-tuning is time-consuming. Additionally, our training pipeline includes diffusion loss, which increases the training time. In cases of extreme changes in video scenes, diffusion loss sometimes fails to guide the model in generating high-quality natural images. These limitations point out the challenge of balancing model adaptability with computational efficiency and effectiveness in varied conditions.

### v2

@@ Abstract
We propose a video editing framework, NaRCan, which integrates a hybrid deformation field and diffusion prior to generate high-quality natural canonical images to represent the input video. Our approach utilizes homography to model global motion and employs multi-layer perceptrons (MLPs) to capture local residual deformations, enhancing the model’s ability to handle complex video dynamics. By introducing a diffusion prior from the early stages of training, our model ensures that the generated images retain a high-quality natural appearance, making the produced canonical images suitable for various downstream tasks in video editing, a capability not achieved by current canonical-based methods. Furthermore, we incorporate low-rank adaptation (LoRA) fine-tuning and introduce a noise and diffusion prior update scheduling technique that accelerates the training process by 14 times. Extensive experimental results show that our method outperforms existing approaches in various video editing tasks and produces coherent and high-quality edited video sequences. Code and video results are available at koi953215.github.io/NaRCan_page.
@@ 1 Introduction
Video editing has always been a fascinating research area. For example, style transfer transforms the original video into a completely new style, enriching the viewing experience. Other tasks include dynamic segmentation and handwriting, which all demonstrate the broad application value of video editing across various fields. Currently, diffusion model technology is becoming increasingly mature and is known for its powerful generative capabilities and frequent use in video editing. However, in video-to-video tasks, maintaining temporal consistency presents a significant challenge, particularly when applying image-based diffusion models to video editing tasks. When these models, originally designed for image generation, are applied frame-by-frame to videos, they often produce temporally inconsistent results due to their frame-independent processing nature. This limitation has motivated numerous research efforts to enhance temporal coherence through various techniques, such as optical flow guidance, latent space alignment, and cross-attention mechanisms, aiming to produce high-quality video sequences while preserving temporal consistency across frames.
Yet, even with solutions addressing temporal consistency, diffusion-based methods [16, 4, 19, 8] still encounter significant challenges in precise localized editing tasks. This is where canonical-based methods demonstrate their unique advantages. By consolidating video content into a single representative image, canonical-based methods [3, 1, 5, 45] enable intuitive and precise spatial control over editing operations while inherently maintaining temporal consistency. This unified representation allows direct application of image-based editing techniques while ensuring that modifications propagate coherently throughout the video sequence, making these methods particularly effective for a wide range of video editing tasks.
While previous canonical-based approaches have established the fundamental utility of canonical images in video editing, our observations suggest that enhancing the naturalness of these representations can significantly improve their effectiveness in various editing scenarios. We observe that existing methods primarily focus on reconstruction quality without explicit constraints to ensure natural canonical image generation. To address this opportunity for enhancement, we propose NaRCan, a novel hybrid deformation field network architecture that incorporates diffusion priors (Figure 1) into the training pipeline (Figure 2). This integration enables our method to generate high-fidelity natural canonical images across diverse scenarios while maintaining temporal consistency. Through comprehensive experimental evaluation, we demonstrate that NaRCan achieves superior performance compared to state-of-the-art video editing methods in both qualitative and quantitative metrics.
@@ 2 Related Work
Some works utilize GAN inversion [65, 78, 60, 74, 48, 26, 69] to edit images or videos. Today, numerous generative models exist for editing images. Some methods, such as GLIDE [44], DALL-E [55, 54], stable diffusion [56], and Imagen [58], are trained on millions of images, resulting in incredible generative abilities. Other methods like SDEdit [39], ControlNet [75], and LDM [2] use conditions to achieve better editing results. Instruction-based video editing methods like InstructPix2Pix [3] and another work [10] often yield sub-optimal results for different editing operations. Techniques like LoRA [18] can assist in fine-tuning to find better weights for editing. Additionally, many zero-shot diffusion-based methods [72, 16, 76, 4, 8, 19, 27] do not require model training but still need constraints to maintain temporal consistency. To address temporal consistency concerns, most methods like Tune-A-Video [70], Text2Video-Zero [27], FateZero [52], and Vid2Vid-Zero [66] incorporate cross-attention mechanisms. Some works propose training for video editing, such as Imagen Video [17] and Make-A-Video [59], but these require large datasets and significant computational resources. Unlike these methods, MeDM [13] uses a flow-coding algorithm to solve this problem. Canonical-based design does not require another mechanism to maintain temporal consistency, however. Once the canonical image is edited, the changes can be propagated to every frame using a deformation field.
@@ 3 Method
In this section, we first introduce our hybrid deformation modeling by combining homography and deformation MLP in Section 3.1. Subsequently, we elaborate on how we integrate the diffusion prior from a LoRA fine-tuned latent diffusion model to ensure the naturalness of our canonical image representation in Section 3.2. Finally, we provide an additional way to improve the quality of our video representation by separating multiple canonical images and describing the necessary changes for downstream tasks in Section 3.3. Figure 2 shows our proposed framework.
@@ 3.1 Hybrid Deformation Field for Deformation Modeling
supplemented with a TVFlow regularization term to prevent overfitting by constraining the magnitude of spatial deformations. While this regularization effectively prevents extreme deformations, it merely imposes magnitude constraints without providing meaningful guidance for modeling complex spatial transformations. To address this limitation, we propose a hybrid deformation field architecture composed of a trainable homography matrix $H(u,v,t)$ and residual deformation MLP:
@@
Unlike the simple constraint-based approach of TVFlow, our method leverages homography to provide global displacement information as structural guidance for the subsequent residual deformation MLP. This hierarchical design enables the residual deformation MLP to learn and express the deformation field more accurately and effectively by building upon the initial geometric transformation provided by the homography matrix.
@@ 4.2 Evaluation
Method Venue Short-term $E_{warp}$downarrow Long-term $E_{warp}$downarrow Interpolation error $\downarrow$
Hashing-nvd [5] ICCV 2023 0.0070 0.0495 9.204
CoDeF [45] CVPR 2024 0.0085 0.0785 8.721
MeDM [13] AAAI 2024 0.0072 0.0583 9.941
Ours - 0.0065 0.0484 8.365
@@
We run our method CoDeF [45], Hashing-nvd [5], CCEdit  [15], and MeDM [13] on the DAVIS [49] and BalanceCC Benchmark that CCEdit proposed for evaluating the results of video editing. In Figure 5, we show the comparison of visual results. The BalanceCC Benchmark provides a unified text prompt for each scene, and we ensure that all video clips from this benchmark are restricted to a maximum of 100 frames. We utilize ControlNet [75] to edit the video using the unified text prompt provided by the BalanceCC Benchmark for a fair comparison. We executed MeDM and CCEdit within their provided environment settings and pre-trained models. Moreover, Hashing-nvd is a video decomposition work that outputs two images representing foreground and background. To maintain the consistent style of these two images for video editing, We utilize the grid trick proposed by RAVE [24] to tackle this problem. Finally, there are also some video-editing results on DAVIS [49], and the corresponding text prompts originated from the BalanceCC Benchmark. For the evaluation of temporal consistency, as shown in Table 1, we utilize short-term [29, 77] and long-term [30] warping errors, along with interpolation error [32], as our primary metrics.
@@
N of separations Training time(s) PSNR $\uparrow$ SSIM $\uparrow$ short-term $E_{warp}$downarrow long-term $E_{warp}$downarrow
1 771.50 23.814 0.6369 0.0016 0.0304
2 1530.79 24.423 0.6762 0.0019 0.0310
3 2275.20 24.852 0.7017 0.0021 0.0312
4 3015.32 25.185 0.7191 0.0022 0.0326
5 3761.40 25.398 0.7301 0.0022 0.0334
Figure 10: Trade-off between reconstruction quality and temporal consistency with varying separations. The visual results demonstrate that increasing the number of separations (from 1 to 5) improves the reconstruction quality of video frames. However, the table reveals a trade-off: as the number of separations increases, temporal consistency decreases. Our empirical findings suggest that using 3 separations typically achieves a balance between reconstruction quality and temporal consistency for most scenarios.
@@ 4.3 Ablation Study
Impact of Separation Numbers.
To investigate the optimal configuration of our Separated NaRCan, we conduct ablation studies by varying the number of canonical images from one to five. Our analysis reveals a clear trade-off between reconstruction quality and temporal consistency, as illustrated in Figure 10. While increasing the number of separations improves frame reconstruction fidelity by allowing more detailed scene representation, it introduces additional transition regions that may compromise temporal coherence. Through extensive experimentation, we find that using three canonical images strikes an optimal balance between representational capacity and temporal stability for most video scenarios.
Figure 11: Failure cases.
@@ 5 Conclusion
Our method relies on LoRA [18] fine-tuning to enhance the diffusion model’s ability to represent the current scene. However, LoRA fine-tuning is time-consuming (about 30 minutes). Additionally, our training pipeline includes diffusion loss, which increases the training time. In cases of extreme changes in video scenes, diffusion loss sometimes fails to guide the model in generating high-quality natural images (Figure 11). These limitations point out the challenge of balancing model adaptability with computational efficiency and effectiveness in varied conditions.
Acknowledgments and Disclosure of Funding
This research was funded by the National Science and Technology Council, Taiwan, under Grants NSTC 112-$2$-E-A49-004-MY2 and 113-2628-E-A49-023-. The authors are grateful to Google, NVIDIA, and MediaTek Inc. for generous donations. Yu-Lun Liu acknowledges the Yushan Young Fellow Program by the MOE in Taiwan.